Circular No. 7982

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V838 MONOCEROTIS S. Desidera and U. Munari, Asiago Astrophysical Observatory, report:

“Spectroscopic observations (ranges 437–774 nm at 0.32 nm/pixel, 625–805 nm at 0.18 nm/pixel, and 819–921 nm at 0.10 nm/pixel) of V838 Mon were obtained on Oct. 2.17 UT with the Asiago 1.82-m telescope (+ AFOSC), following conjunction with the sun. Longward of 700 nm, the spectrum is dominated by TiO and particularly VO absorption bands, suggesting a spectral type slightly later than the M10-III star U Her. Thus, a dramatic temperature decrease of the outbursting star is contributing to the fading at optical wavelengths and to the brightening in the infrared observed over the last few months. We see no Hα in emission, no Ca II triplet absorption in the far red, and no P-Cyg profiles elsewhere, thus marking a sharp contrast with the last spectra recorded in April–May. Shortward of 700 nm, a weak blue continuum dominates the spectrum. At the low S/N of our observations, the spectrum could resemble the F0 type estimated in quiescence for V838 Mon by Munari et al. (2002, A.Ap. 389, L51), in which case the eruption that took V838 Mon to V = 6.7 in Feb. could have occurred in a binary companion to the F0 star. Deeper observations of the blue continuum are in progress.”

SUPERNOVA 2001gd IN NGC 5033 P. Chandra and A. Ray, Tata Institute of Fundamental Research,

Mumbai; and S. Bhatnagar, National Radio Astronomy Observatory, Socorro, report the detection of SN 2001gd at 1396.75 MHz (flux density 3.55 ± 0.22 mJy) with the Giant Meterwave Radio Telescope (GMRT) on Sept. 22.22 UT. The extended emission from the host galaxy had a flux density of 16.2 ± 1.7 mJy in a 8′′.6 × 6′′.8 box centered at α = 13^{h}13^{m}27^{s}.5, δ = +36^{o}35′36′′ (equinox 2000.0). SN 2001gd was not detected (2σ upper limit of 0.50 mJy) with the GMRT at 616.75 MHz on May 19.84. The flux from the host-galaxy’s center at this frequency was 92 ± 7.2 mJy in a 29′′.9 × 28′′.2 box centered at α = 13^{h}13^{m}27^{s}.4, δ = +36^{o}35′39′′.

V4743 SAGITTARII N. Brown, Quinns Rocks, W. Australia, reports that this nova was not

present on nova patrol films taken on Sept. 9.6 UT (limiting mag 12.0). Visual magnitude estimates by A. Pearce, Nedlands, W. Australia: Sept. 25.505, 6.5; 26.486, 6.8; 27.498, 7.1; 29.480, 8.1; Oct. 2.535, 8.1.

2002 October 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green